# Beyond the EPICS: comprehensive Python IOC development with QueueIOC

### v3

@@ 7 Conclusion
Architectural deficiencies in EPICS lead to inefficiency in development and application; from the perspective of complexity and succinctness, the essence of these problems is that the architecture of EPICS makes the complexities of EPICS IOCs vastly higher than the lower bounds. A backward-compatible way to avoid these problems is replacing EPICS IOCs with Python IOCs based on libraries like caproto. Learning from EPICS OPIs, we can require widgets in GUI frontends to communicate only with the main event loop, and mandate the use of message passing for this communication; based on this idea, the submit/notify pattern is formed, which is also related to the actor/CSP models and the MVC pattern. Mamba frontends and standalone GUI frontends following the pattern have been developed; the communication between both kinds of frontends and their backends may be categorised into requests/replies and notifications, which are handled in main event loops inside the backends. The combination of an event loop, requests and notifications can also be observed elsewhere; thus by treating caput, caget and camonitor as specialised requests/replies and notifications, and handling them in Python-based main loops, the QueueIOC framework is formed. After comparing the architecture of QueueIOC with that of EPICS, we believe QueueIOC has the potential to eventually replace most EPICS IOCs currently used with succinct workalikes; under certain conditions, it may even become a starting point for a unified device-control ecosystem. Examples given for QueueIOC include workalikes of StreamDevice/asyn, as well as seq-like IOCs for monochromators, motor anti-bumping and motor multiplexing; also reported are a workalike of procServ, as well as a procServControl workalike based on it and QueueIOC. A QueueIOC-based framework for detector integration, which aims to overcome some architectural limitations of areaDetector while still offering decent performance, is presented in (Zhang et al., 2024). A practical analysis is given for the unique benefits of QueueIOC, emphasising its pursuit of the utmost simplicity, which leads to significant reduction in the cost of development and maintenance.
Acknowledgements
@@
Liu would like to thank the Linux Club of Peking University for providing a venue in 2019 to share his preliminary thoughts on replacing EPICS (https://lcpu.club/wiki/index.php?title=2019%E6%B4%BB%E5%8A%A8B07). Much more importantly, LCPU introduced him to the wonderland of Unix and Lisp, where the inspirations for this paper came from; finding solutions much simpler than what have been previously imagined, just like those in (Sussman and Steele, 1998) and (Landley, 2017), has become a continuous source of joy and motivation. This work was supported by the National Key Research and Development Program for Young Scientists (Grant No. 2023YFA1609900) and the Young Scientists Fund of the National Natural Science Foundation of China (Grants Nos. 12205328, 12305371).

### v4

@@ 7 Conclusion
Architectural deficiencies in EPICS lead to inefficiency in development and application; from the perspective of complexity and succinctness, the essence of these problems is that the architecture of EPICS makes the complexities of EPICS IOCs vastly higher than the lower bounds. A backward-compatible way to avoid these problems is replacing EPICS IOCs with Python IOCs based on libraries like caproto. Learning from EPICS OPIs, we can require widgets in GUI frontends to communicate only with the main event loop, and mandate the use of message passing for this communication; based on this idea, the submit/notify pattern is formed, which is also related to the actor/CSP models and the MVC pattern. Mamba frontends and standalone GUI frontends following the pattern have been developed; the communication between both kinds of frontends and their backends may be categorised into requests/replies and notifications, which are handled in main event loops inside the backends. The combination of an event loop, requests and notifications can also be observed elsewhere; thus by treating caput, caget and camonitor as specialised requests/replies and notifications, and handling them in Python-based main loops, the QueueIOC framework is formed. After comparing the architecture of QueueIOC with that of EPICS, we believe QueueIOC has the potential to eventually replace most EPICS IOCs currently used with succinct workalikes; under certain conditions, it may even become a starting point for a unified device-control ecosystem. Examples given for QueueIOC include workalikes of StreamDevice/asyn, as well as “sequencer” IOCs (like those based on seq), including those for monochromators, motor anti-bumping and motor multiplexing; also reported are a workalike of procServ, as well as a procServControl workalike based on it and QueueIOC. A QueueIOC-based framework for detector integration, which aims to overcome some architectural limitations of areaDetector while still offering decent performance, is presented in [19]. A practical analysis is given for the unique benefits of QueueIOC, emphasising its pursuit of the utmost simplicity, which leads to significant reduction in the cost of development and maintenance.
Statements and declarations
Acknowledgements:
@@
Liu would like to thank the Linux Club of Peking University for providing a venue in 2019 to share his preliminary thoughts on replacing EPICS (https://lcpu.club/wiki/index.php?title=2019%E6%B4%BB%E5%8A%A8B07). Much more importantly, LCPU introduced him to the wonderland of Unix and Lisp, where the inspirations for this paper came from; finding solutions much simpler than what have been previously imagined, just like those in [22] and [23], has become a continuous source of joy and motivation.
Funding:
This work was supported by the National Key Research and Development Program for Young Scientists (Grant No. 2023YFA1609900) and the Young Scientists Fund of the National Natural Science Foundation of China (Grants Nos. 12205328, 12305371).
Data availability:
QueueIOC and s6-epics have been released, respectively, at https://codeberg.org/CasperVector/queue_iocs and https://codeberg.org/CasperVector/s6-epics; fully open-source editions of Mamba and ihep-pkg have been released, respectively, at https://codeberg.org/CasperVector/mamba-ose and https://codeberg.org/CasperVector/ihep-pkg-ose. QueueIOC itself depends on some patches for caproto; GUIs in the open-source edition of Mamba (including the standalone ones, which are in the mamba_lite subdirectory) depend on some patches for pyqtgraph; the supplementary materials include OPIs which depend on some patches for PyDM; all these currently HEPS-specific patches are available from the open-source edition of ihep-pkg.